OpenStudy (anonymous):

Write the function y=3(x+b)^2-7 in standard form

OpenStudy (anonymous):

\[3(x+b)^2-7=3(x+b)(a+b)-7=3(x^2+2xb+b^2)-7\] \[=3x^2+6xb+3b^2-7\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Just expansion.

OpenStudy (anonymous):

btw how did yu get the 2xb

OpenStudy (anonymous):

Use the formula ( a + b) whole square. Here, (x + b) ^ 2= x^2 + b^2 + 2xb